# Geometric Constraints on Page Curves: Insights from Island Rule and Quantum Focusing Conjecture

### v1

@@ 1 Introduction
Recently, studies have demonstrated that the island formula does not depend on the AdS/CFT correspondence and has been applied far beyond the asymptotically AdS black holes. One can refer to a non-exhaustive list of progress in this field eternalbh ; jt1 ; jt2 ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; 2d7 ; 2d8 ; 2d9 ; high1 ; high2 ; high3 ; high4 ; high5 ; high6 ; high7 ; high8 ; high9 ; high10 ; high11 ; high12 ; high13 ; high14 ; high15 ; high16 ; high17 ; high18 ; high19 ; high20 ; high21 ; high22 ; high23 ; high24 ; high25 ; high26 ; high27 ; high28 ; high29 ; high30 ; high31 ; high32 ; high33 ; high34 ; high35 ; high36 ; high37 ; high38 ; high39 ; high40 ; high41 ; high42 ; high43 ; high44 ; high45 ; high46 ; high47 ; high48 ; high49 ; high50 .
@@ 4 Discussion and Conclusion
In conclusion, our study significantly contributes to the comprehension of black hole evaporation dynamics and the resolution of the information paradox, leveraging the insights from the Page curve and island paradigm. By examining diverse static black hole metrics within asymptotically flat or (anti-)de Sitter spacetimes, we revealed a pivotal geometric requirement: A metric with a negative second derivative near the horizon, $f^{\prime\prime}(r)<0$, in proximity to the event horizon where $r\sim r_{h}+{\cal O}(G_{N})$, which is fundamental for island formation and the consequent derivation of Page curves. This discovery affirms the universality of Page curves, transcending model-specific restrictions and reinforcing the compatibility of information conservation within the semi-classical gravity framework.
@@ Acknowledgements.
We would thank to Shuyi Lin, Ruidong Zhu and Xiaokai He for helpful discussions. The study was partially supported by NSFC, China (Grant No. 12275166 and No. 12311540141).

### v2

# Geometric Constraints via Page Curves: Insights from Island Rule and Quantum Focusing Conjecture

@@ 1 Introduction
Recently, studies have demonstrated that the island formula does not depend on the AdS/CFT correspondence and has been applied far beyond the asymptotically AdS black holes. One can refer to a non-exhaustive list of progress in this field eternalbh ; jt1 ; jt2 ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; 2d7 ; 2d8 ; 2d9 ; high1 ; high2 ; high3 ; high4 ; high5 ; high6 ; high7 ; high8 ; high9 ; high10 ; high11 ; high12 ; high13 ; high14 ; high15 ; high16 ; high17 ; high18 ; high19 ; high20 ; high21 ; high22 ; high23 ; high24 ; high25 ; high26 ; high27 ; high28 ; high29 ; high30 ; high31 ; high32 ; high33 ; high34 ; high35 ; high36 ; high37 ; high38 ; high39 ; high40 ; high41 ; high42 ; high43 ; high44 ; high45 ; high46 ; high47 ; high48 ; high49 ; high50 ; high51 ; high52 ; high53 ; high54 ; high55 ; high56 ; high57 .
@@ 4 Discussion and Conclusion
In conclusion, our study significantly contributes to the comprehension of black hole evaporation dynamics and the resolution of the information paradox, leveraging the insights from the Page curve and island paradigm. By examining diverse static black hole metrics within asymptotically flat or (anti-)de Sitter spacetimes, we revealed a pivotal geometric requirement: A metric with a negative second derivative near the horizon, $f^{\prime\prime}(r)<0$, in proximity to the event horizon where $r\sim r_{h}+{\cal O}(G_{N})$. which is fundamental for island formation and the consequent derivation of Page curves. This discovery affirms the universality of Page curves, transcending model-specific restrictions and reinforcing the compatibility of information conservation within the semi-classical gravity framework.
@@ Acknowledgements.
We would thank to Shuyi Lin, Ruidong Zhu and Xiaokai He for discussions related to island rule. The study was partially supported by NSFC, China (Grant No. 12275166 and No. 12311540141).